# The Bisognano-Wichmann property for non-unitary Wightman conformal field theories

### v1

@@ Abstract
The Bisognano-Wichmann and Haag duality properties for algebraic quantum field theories are often studied using the powerful tools of Tomita-Takesaki modular theory for nets of operator algebras. In this article, we study analogous properties of nets of algebras generated by smeared Wightman fields, for potentially non-unitary theories. In light of recent work [CRTT24] constructing Wightman field theories for (non-unitary) Möbius vertex algebras, we obtain a broadly applicable non-unitary version of the Bisognano-Wichmann property. In this setting we do not have access to the traditional tools of Hilbert space functional analysis, like functional calculus. Instead, results analogous to those of Tomita-Takesaki theory are derived ‘by hand’ from the Wightman axioms. As an application, we demonstrate Haag duality for nets of smeared Wightman fields for unitary theories.
@@ 1 Introduction
It has long been understood that the axioms of a vertex algebra (the most common way of axiomatizing a non-unitary conformal field theory) are closely tied to the Wightman axioms [Kac98, §1.2], and in recent joint work we showed that there is an equivalence of categories between (potentially non-unitary) Möbius-covariant Wightman CFTs on the unit circle $S^{1}$ and (potentially non-unitary) Möbius vertex algebras which are generated by a chosen set of quasiprimary vectors [CRTT24, RTT22]. Such a Wightman theory consists of a family ${\mathcal{F}}$ of operator-valued distributions $\varphi:C^{\infty}(S^{1})\to\operatorname{End}({\mathcal{D}})$ on a common, invariant domain ${\mathcal{D}}$ cyclically generated from a vacuum vector $\Omega$, along with a compatible positive-energy representation $U$ of the Möbius group $\operatorname{M\ddot{o}b}=\operatorname{PSU}(1,1)$ of holomorphic automorphisms of the unit disk.
Following [SW64], we may then form for each interval111i.e. open connected nonempty non-dense subset $I\subset S^{1}$ the polynomial algebra ${\mathcal{P}}(I)$ generated by all smeared fields $\varphi(f)$, where $\varphi\in{\mathcal{F}}$ and $f\in C^{\infty}(S^{1})$ is supported in $I$. These polynomial algebras formally satisfy many of the same axioms as a Haag-Kastler net of operator algebras, but they are actually built from a vertex algebra. Our philosophy therefore is to try to obtain analogous results for the polynomial nets ${\mathcal{P}}$ as one does with Haag-Kastler nets, and in doing so apply operator algebraic methods in the non-unitary setting. An early success in this regard was the proof of the Reeh-Schlieder property that $\Omega$ is cyclic and separating for the algebras ${\mathcal{P}}(I)$ for non-unitary Wightman theories [CRTT24, App. A].
In practice, transferring operator algebraic methods to the non-unitary setting can be quite difficult as these methods often rely crucially on features of operators on Hilbert spaces, e.g. by applying functional calculus to produce operators with certain properties. However, in this article we report success in this endeavor with regard to the Bisognano-Wichmann property, and partial success with regard to Haag duality. We are guided by results proven in the setting of unitary CFTs, but often a novel approach is required due to the lack of general functional analytic theory. For example, one cannot invoke the general Tomita-Takesaki theory, but instead the key results can be established ‘by hand’ through careful study of the CFT axioms; the same is true with regard to applications of functional calculus.
@@
be the one-parameter subgroup of $\operatorname{M\ddot{o}b}$ leaving the upper- and lower-semicircles $I_{\pm}$ invariant (which uniquely determines $v_{t}$ up to rescaling the parameter $t$), and let $V_{t}=U(v_{t})\in\operatorname{End}({\mathcal{D}})$. The analytic continuations $\widetilde{V}_{\pm i\pi}$ have domains $D(\widetilde{V}_{\pm i\pi})\subset{\mathcal{D}}$ given by vectors $\Phi\in{\mathcal{D}}$ for which there exists a $\Psi\in{\mathcal{D}}$ such that the following holds: Let ${\mathbb{S}}_{\pm\pi i}\subset{\mathbb{C}}$ be the closed strip bounded by the lines ${\mathbb{R}}$ and ±${\mathbb{R}}\pm\pi i$. For every continuous (with respect to the canonical ${\mathcal{F}}$-strong topology on ${\mathcal{D}}$) linear functional $\lambda:{\mathcal{D}}\to{\mathbb{C}}$, there exists a function $G_{\lambda}:{\mathbb{S}}_{\pm\pi i}\to{\mathbb{C}}$, continuous on the closed strip and holomorphic on the interior, such that
@@
The Bisognano-Wichmann property for a non-unitary CFT with no extra structure is the following (proven as Theorem 3.10).
@@
It is not too difficult to establish the result of Theorem A when the functions $f_{j}$ lie in a dense subspace, but in the absence of the usual Hilbert space tools, extending this result to all of ${\mathcal{P}}(I_{\pm})\Omega$ requires a careful study of the analytic functions (1.1). We do not assume anything like finite-dimensionality of the conformal weight spaces, and so Theorem A applies to a variety of interesting log CFTs, such as the $\beta\gamma$-ghost system with $c=2$ (typically studied as a vertex algebra, but realizable as a Wightman CFT by [CRTT24]).
[…]
@@
in the strong sense of partially defined operators on ${\mathcal{D}}$ (Theorem 4.6).
@@
The above is the familiar Bisognano-Wichmann property linking the analytic continuation of $V_{t}$, the PCT operator, and the adjoint operation; it appears as Theorem 4.2. We also obtain the KMS property as Corollary 4.3:
@@
As in the motivating articles [BW75, BSM90, BGL93, GF93], it is natural to next turn to questions of Haag duality. Let $\operatorname{End}_{*}({\mathcal{D}})$ be the *-algebra of continuous endomorphisms of ${\mathcal{D}}$ with a continuous adjoint. Let ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$, where the notion $S^{\prime}$ refers to the commutant in $\operatorname{End}_{*}({\mathcal{D}})$.
Conjecture D.
@@
Here $I^{\prime}$ denotes the (interior of the) complementary interval. The importance of Conjecture D lies in the ubiquitous role of Haag duality in the analysis of Haag-Kastler nets of operator algebras, particularly with regard to their representation theory. A positive resolution of the conjecture would pave the way for further analysis of the representation theory of non-unitary Wightman polynomial nets mimicking operator algebraic methods. We obtain several sufficient conditions for Conjecture D in Proposition 5.18. One such condition is that ${\mathcal{P}}(I_{+})\Omega$ is a core for $V_{i\pi}$, in which case the result of Theorem B can be written as an operator identity of the form “$S=\theta V_{i\pi}$” familiar from Tomita-Takesaki theory and unitary Bisognano-Wichmann phenomena.
We do, however, provide a positive answer to Conjecture D when the Wightman CFT is unitary, in Corollary 6.2.
@@
Let $({\mathcal{F}},{\mathcal{D}},U,\Omega)$ be a unitary Wightman CFT on $S^{1}$. Let ${\mathcal{P}}(I)\subset\operatorname{End}_{*}({\mathcal{D}})$ be the unital algebra generated by smeared fields $\varphi(f)$ with $f$ supported in $I$, and let ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$ with the commutant taken in $\operatorname{End}_{*}({\mathcal{D}})$. Then ${\mathcal{Q}}(I^{\prime})={\mathcal{Q}}(I)^{\prime}$.
Let ${\mathcal{P}}(I)_{sa}$ denote the real subspace of self-adjoint elements of ${\mathcal{P}}(I)$. The proof of Theorem E relies on showing that the norm-closed real subspaces $\overline{{\mathcal{P}}(I)_{sa}\Omega}$ of the Hilbert space completion of ${\mathcal{D}}$ form a net of standard subspaces in the sense of [Lon08a]. Even in the unitary case, this result provides a new avenue to apply operator algebraic methods to study the representation theory of unitary vertex operator algebras; we hope to take up this idea in future work.
Theorem F.
Let ${\mathcal{V}}$ be a unitary Möbius vertex algebra with vacuum vector $\Omega$, and let ${\mathcal{A}}(I)$ be the net of von Neumann algebras described above. Then ${\mathcal{V}}$ is AQFT-local if and only if $\Omega$ is a separating vector for ${\mathcal{A}}(I)$.
@@
This article is dedicated to Huzihiro Araki, and was prepared for the special issue of Communications in Mathematical Physics in his honor. We briefly explain the connection between this article and Araki’s work. Araki’s proof of Haag duality for a free scalar bosonic field [Ara63, Ara64], sometimes called Haag-Araki duality, laid the foundation for later study of duality using the methods of Bisognano-Wichmann; this article is built on that foundation. Araki also considered duality for free fermionic fields [Ara71, Rem. 4.9], and this instance of Haag-Araki duality plays a special role with regard to the philosophy of this article. Indeed, we are motivated here by the comparison between nets ${\mathcal{P}}(I)$ generated by (typically unbounded) smeared fields, and the corresponding nets of von Neumann algebras ${\mathcal{A}}(I)$. The free fermion model is essentially the only example of a 2d chiral CFT whose smeared fields are bounded operators, so that the distinction between ${\mathcal{P}}(I)$ and ${\mathcal{A}}(I)$ is not necessary. Our Theorem E establishing Haag duality for nets of smeared fields can therefore be interpreted as a generalization of Haag-Araki duality to non-free models.
@@ Acknowledgements
The author was supported by ARC Discovery Project DP200100067.
@@ 2.1 Conformal nets and Tomita-Takesaki theory
Our first aim is to describe the relationship between Tomita-Takesaki modular theory and algebraic conformal field theories on Hilbert spaces which provides the motivation for this work. We begin by summarizing the main results of Tomita-Takesaki theory, which was first introduced [Tom67, Tak70] in the context of von Neumann algebras (see also [Tak03]). If ${\mathcal{H}}$ is a Hilbert space, then a von Neumann algebra $M$ is a unital *-subalgebra $M\subset{\mathcal{B}}({\mathcal{H}})$ that is closed in the topology of pointwise convergence. A vector $\Omega\in{\mathcal{H}}$ is called cyclic for $M$ if $M\Omega$ is dense in ${\mathcal{H}}$, and called separating if the only element $x\in M$ satisfying $x\Omega=0$ is $x=0$. A vector $\Omega$ is cyclic for $M$ if and only if it is separating for $M^{\prime}$, and in light of the double commutant theorem $M=M^{\prime\prime}$, it follows that $\Omega$ is cyclic for $M^{\prime}$ if and only if it is separating for $M$ [^3] Here the commutant $S^{\prime}$ of a subset $S\subset{\mathcal{B}}({\mathcal{H}})$ consists of all operators $x\in{\mathcal{B}}({\mathcal{H}})$ commuting element-wise with $S$. This is a von Neumann algebra if $S$ is closed under taking adjoints.
Let $S^{0}$ be the densely-defined conjugate linear operator on $H$ which maps $M\Omega$ to itself via the formula $S^{0}x\Omega=x^{*}\Omega$, and let $S$ be the closure of $S^{0}$ (i.e. the operator obtained by taking the closure of the graph of $S^{0}$ in ${\mathcal{H}}\times{\mathcal{H}}$). We have a polar decomposition $S=J\Delta^{1/2}$, where $J$ is antiunitary and $\Delta^{1/2}$ is a densely defined positive operator. The main result of Tomita-Takesaki theory says that
@@
Tomita-Takesaki theory has proven to be an extremely powerful tool in the study of algebraic quantum field, which is to say quantum field theories in the sense of Haag-Kastler nets of operator algebras [HK64]. In the context of 2d chiral conformal field theory, we have the following axiomatization:
@@
${\mathcal{K}}^{\prime}=\{\xi\in{\mathcal{H}}\,:\,\langle\xi,\eta\rangle\in{\mathbb{R}}\,\text{for all}\,\eta\in{\mathcal{K}}.\}$
Alternatively, ${\mathcal{K}}^{\prime}=(i{\mathcal{K}})^{\perp}$ where the orthogonal complement is taken with respect to the real Hilbert space inner product $\operatorname{Re}\langle\,\cdot\,,\,\cdot\,\rangle$ (from which it follows that ${\mathcal{K}}^{\prime\prime}={\mathcal{K}}$). Following [Lon08a], one has the following notion of a net of standard subspaces:

### v2

@@ Abstract
The Bisognano-Wichmann and Haag duality properties for algebraic quantum field theories are often studied using the powerful tools of Tomita-Takesaki modular theory for nets of operator algebras. In this article, we study analogous properties of nets of algebras generated by smeared Wightman fields, for potentially non-unitary theories. In light of recent work [CRTT25] constructing Wightman field theories for (non-unitary) Möbius vertex algebras, we obtain a broadly applicable non-unitary version of the Bisognano-Wichmann property. In this setting we do not have access to the traditional tools of Hilbert space functional analysis, like functional calculus. Instead, results analogous to those of Tomita-Takesaki theory are derived ‘by hand’ from the Wightman axioms. As an application, we demonstrate Haag duality for nets of smeared Wightman fields.
@@ 1 Introduction
It has long been understood that the axioms of a vertex algebra (the most common way of axiomatizing a non-unitary conformal field theory) are closely tied to the Wightman axioms [Kac98, §1.2], and in recent joint work we showed that there is an equivalence of categories between (potentially non-unitary) Möbius-covariant Wightman CFTs on the unit circle $S^{1}$ and (potentially non-unitary) Möbius vertex algebras which are generated by a chosen set of quasiprimary vectors [CRTT25, RTT22]. Such a Wightman theory consists of a family $\mathcal{F}$} of operator-valued distributions $\varphi:C^{\infty}(S^{1})\to\operatorname{End}({\mathcal{D}})$ on a common, invariant domain $\mathcal{D}$} cyclically generated from a vacuum vector $\Omega$, along with a compatible positive-energy representation $U$ of the Möbius group $\operatorname{M\ddot{o}b}=\operatorname{PSU}(1,1)$ of holomorphic automorphisms of the unit disk.
Following [SW64], we may then form for each interval111i.e. open connected nonempty non-dense subset $I\subset S^{1}$ the polynomial algebra ${\mathcal{P}}(I)$ generated by all smeared fields $\varphi(f)$, where $\varphi\in{\mathcal{F}}$ and $f\in C^{\infty}(S^{1})$ is supported in $I$. These polynomial algebras formally satisfy many of the same axioms as a Haag-Kastler net of operator algebras, but they are actually built from a vertex algebra. Our philosophy therefore is to try to obtain analogous results for the polynomial nets $\mathcal{P}$} as one does with Haag-Kastler nets, and in doing so apply operator algebraic methods in the non-unitary setting. An early success in this regard was the proof of the Reeh-Schlieder property that $\Omega$ is cyclic and separating for the algebras ${\mathcal{P}}(I)$ for non-unitary Wightman theories [CRTT25, App. A].
In practice, transferring operator algebraic methods to the non-unitary setting can be quite difficult as these methods often rely crucially on features of operators on Hilbert spaces, e.g. by applying functional calculus to produce operators with certain properties. However, in this article we report success in this endeavor with regard to the Bisognano-Wichmann property and Haag duality. We are guided by results proven in the setting of unitary CFTs, but often a novel approach is required due to the lack of general functional analytic theory. For example, one cannot invoke the general Tomita-Takesaki theory, but instead the key results can be established ‘by hand’ through careful study of the CFT axioms; the same is true with regard to applications of functional calculus.
@@
be the one-parameter subgroup of $\operatorname{M\ddot{o}b}$ leaving the upper- and lower-semicircles $I_{\pm}$ invariant (which uniquely determines $v_{t}$ up to rescaling the parameter $t$), and let $V_{t}=U(v_{t})\in\operatorname{End}({\mathcal{D}})$. The analytic continuations $\widetilde{V}_{\pm i\pi}$ have domains $D(\widetilde{V}_{\pm i\pi})\subset{\mathcal{D}}$ given by vectors $\Phi\in{\mathcal{D}}$ for which there exists a $\Psi\in{\mathcal{D}}$ such that the following holds: Let ${\mathbb{S}}_{\pm\pi i}\subset{\mathbb{C}}$ be the closed strip bounded by the lines $\mathbb{R}$} and ${\mathbb{R}}\pm\pi i$. For every continuous (with respect to the canonical $\mathcal{F}$}-strong topology on $\mathcal{D}$}) linear functional $\lambda:{\mathcal{D}}\to{\mathbb{C}}$, there exists a function $G_{\lambda}:{\mathbb{S}}_{\pm\pi i}\to{\mathbb{C}}$, continuous on the closed strip and holomorphic on the interior, such that
@@
Traditionally, the Bisognano-Wichmann property has been formulated in a unitary setting, identifying the analytic continuation of the group $V_{t}$ with the PCT operator and the adjoint operation on smeared fields. On the other hand, for a non-unitary CFT with no extra structure, one may not have a PCT operator or adjoint operation. Even in this general setting, however, we are still able to compute the analytic continuation of the group $V_{t}$ (Theorem 3.10 in the main body of the article):
@@
It is not too difficult to establish the result of Theorem A when the functions $f_{j}$ lie in a dense subspace, but in the absence of the usual Hilbert space tools, extending this result to all of ${\mathcal{P}}(I_{\pm})\Omega$ requires a careful study of the analytic functions (1.1). We do not assume anything like finite-dimensionality of the conformal weight spaces, and so Theorem A applies to a variety of interesting log CFTs, such as the $\beta\gamma$-ghost system with $c=2$ (typically studied as a vertex algebra, but realizable as a Wightman CFT by [CRTT25]).
In order to say more, we will now assume that our Wightman CFT possesses an invariant nondegenerate Hermitian form $\langle\,\cdot\,$$\cdot\,\rangle$ on $\mathcal{D}$}, along with its antiunitary PCT operator $\theta$. Such a theory would be unitary if we assumed the form to be positive definite, but there are important non-unitary examples motivating our study, such as the non-unitary minimal models (e.g. the Yang-Lee model). In this case we have a notion of adjoint of an operator on $\mathcal{D}$} (although the adjoint is not guaranteed to exist), and the polynomial algebras become *-algebras. In this setting, we consider a slightly modified version of the operators $\widetilde{V}_{\pm i\pi}$, which we denote $V_{\pm i\pi}$, for which we only require analytic continuations (1.1) to exist for linear functionals $\lambda$ of the form $\langle\,\cdot\,,\Phi\rangle$. These modified operators are self-adjoint:
@@
in the strong sense of partially defined operators on $\mathcal{D}$} (Theorem 4.6). For these operators we obtain the familiar Bisognano-Wichmann property linking the analytic continuation of $V_{t}$, the PCT operator, and the adjoint operation; it appears as Theorem 4.2 in the body of the text.
@@
We also obtain the KMS property as Corollary 4.3:
@@
As in the motivating articles [BW75, BSM90, BGL93, GF93], it is natural to next turn to questions of Haag duality. Let $\operatorname{End}_{*}({\mathcal{D}})$ be the *-algebra of continuous endomorphisms of $\mathcal{D}$} with a continuous adjoint. The algebras ${\mathcal{P}}(I)$ are a non-unitary Haag-Kastler nets of subalgebras of $\operatorname{End}_{*}({\mathcal{D}})$, as defined in Section 5. This means that they satisfy: (i) isotony - ${\mathcal{P}}(I)\subset{\mathcal{P}}(J)$ when $I\subset J$, (ii) Möbius covariance - they transform appropriately under a representation of $\operatorname{M\ddot{o}b}$, (iii) locality - ${\mathcal{P}}(I)$ and ${\mathcal{P}}(J)$ commute when $I$ and $J$ are disjoint, and (iv) vacuum - there is a chosen vector $\Omega$ which is cyclic for each algebra ${\mathcal{P}}(I)$. Let ${\mathcal{Q}}(I)={\mathcal{P}}(I)^{\prime\prime}$, where the notion $S^{\prime}$ refers to the commutant in $\operatorname{End}_{*}({\mathcal{D}})$. Then ${\mathcal{Q}}(I)$ is again a Haag-Kastler net of subalgebras of $\operatorname{End}_{*}({\mathcal{D}})$, and ${\mathcal{P}}(I)\subseteq{\mathcal{Q}}(I)$ for all intervals $I$. The Haag duality property of $\mathcal{Q}$} asserts that ${\mathcal{Q}}(I^{\prime})={\mathcal{Q}}(I)^{\prime}$, where $I^{\prime}$ denotes the (interior of the) complementary interval. We establish Haag duality in Theorem 5.18.
Theorem D.
@@
The importance of Theorem D lies in the ubiquitous role of Haag duality in the analysis of Haag-Kastler nets of operator algebras, particularly with regard to their representation theory. In this context, the operation of ‘double commutant’ is quite natural, as it provides the von Neumann algebra generated by a unital *-subalgebra of the bounded operators on a Hilbert space. We hope that Theorem D opens up new avenues for applying Haag duality to the representation theory of non-unitary Wightman polynomial nets mimicking operator algebraic methods.
Let ${\mathcal{P}}(I)_{sa}$ denote the real subspace of self-adjoint elements of ${\mathcal{P}}(I)$. The proof of Theorem D relies on studying the family of $\mathcal{F}$}-strongly closed real subspaces ${\mathcal{K}}(I):=\overline{{\mathcal{P}}(I)_{sa}\Omega}\subseteq{\mathcal{D}}$. These are a non-unitary analog of a net of standard subspaces in the sense of [Lon08a]. Even in the unitary case, this result provides new avenues to apply operator algebraic methods to study unitary vertex operator algebras.
@@
Let $\mathcal{V}$} be a unitary Möbius vertex algebra with vacuum vector $\Omega$, and let ${\mathcal{A}}(I)$ be the net of von Neumann algebras described above. Then $\mathcal{V}$} is AQFT-local if and only if for some interval $I$ the vector $\Omega$ is a separating vector for ${\mathcal{A}}(I)$.
@@
This article is dedicated to Huzihiro Araki, and was prepared for the special issue of Communications in Mathematical Physics in his honor. We briefly explain the connection between this article and Araki’s work. Araki’s proof of Haag duality for a free scalar bosonic field [Ara63, Ara64], sometimes called Haag-Araki duality, laid the foundation for later study of duality using the methods of Bisognano-Wichmann; this article is built on that foundation. Araki also considered duality for free fermionic fields [Ara71, Rem. 4.9], and this instance of Haag-Araki duality plays a special role with regard to the philosophy of this article. Indeed, we are motivated here by the comparison between nets ${\mathcal{P}}(I)$ generated by (typically unbounded) smeared fields, and the corresponding nets of von Neumann algebras ${\mathcal{A}}(I)$. The free fermion model is essentially the only example of a 2d chiral CFT whose smeared fields are bounded operators, so that the distinction between ${\mathcal{P}}(I)$ and ${\mathcal{A}}(I)$ is not necessary. Our Theorem D establishing Haag duality for nets of smeared fields can therefore be interpreted as a generalization of Haag-Araki duality to non-free models.
@@ Acknowledgements
The author was supported by ARC Discovery Project DP200100067. The author would like to thank the anonymous referee for helpful suggestions.
@@ 2.1 Conformal nets and Tomita-Takesaki theory
Our first aim is to describe the relationship between Tomita-Takesaki modular theory and algebraic conformal field theories on Hilbert spaces which provides the motivation for this work. We begin by summarizing the main results of Tomita-Takesaki theory, which was first introduced [Tom67, Tak70] in the context of von Neumann algebras (see also [Tak03]). If $\mathcal{H}$} is a Hilbert space, then a von Neumann algebra $M$ is a unital *-subalgebra $M\subset{\mathcal{B}}({\mathcal{H}})$ that is closed in the topology of pointwise convergence. A vector $\Omega\in{\mathcal{H}}$ is called cyclic for $M$ if $M\Omega$ is dense in $\mathcal{H}$}, and called separating if the only element $x\in M$ satisfying $x\Omega=0$ is $x=0$. A vector $\Omega$ is cyclic for $M$ if and only if it is separating for $M^{\prime}$, and in light of the double commutant theorem $M=M^{\prime\prime}$, it follows that $\Omega$ is cyclic for $M^{\prime}$ if and only if it is separating for MM [^3] Here the commutant $S^{\prime}$ of a subset $S\subset{\mathcal{B}}({\mathcal{H}})$ consists of all operators $x\in{\mathcal{B}}({\mathcal{H}})$ commuting element-wise with $S$. This is a von Neumann algebra if $S$ is closed under taking adjoints.
Let $S^{0}$ be the densely-defined conjugate linear operator on $\mathcal{H}$} which maps $M\Omega$ to itself via the formula $S^{0}x\Omega=x^{*}\Omega$, and let $S$ be the closure of $S^{0}$ (i.e. the operator obtained by taking the closure of the graph of $S^{0}$ in $\mathcal{H}$$\times{\mathcal{H}}$). We have a polar decomposition $S=J\Delta^{1/2}$, where $J$ is antiunitary and $\Delta^{1/2}$ is a densely defined positive operator. The main result of Tomita-Takesaki theory says that
@@
Tomita-Takesaki theory has proven to be an extremely powerful tool in the study of algebraic quantum field theories, which is to say quantum field theories in the sense of Haag-Kastler nets of operator algebras [HK64]. In the context of 2d chiral conformal field theory, we have the following axiomatization:
@@
${\mathcal{K}}^{\prime}=\{\xi\in{\mathcal{H}}\,:\,\langle\xi,\eta\rangle\in{\mathbb{R}}\,\text{ for all }\,\eta\in{\mathcal{K}}\}.$
Alternatively, ${\mathcal{K}}^{\prime}=(i{\mathcal{K}})^{\perp}$ where the orthogonal complement is taken with respect to the real Hilbert space inner product $\operatorname{Re}\langle\,\cdot\,$$\cdot\,\rangle$ (from which it follows that ${\mathcal{K}}^{\prime\prime}={\mathcal{K}}$). Following [Lon08a], one has the following notion of a net of standard subspaces: